OpenStudy (anonymous):

2radical7.6+2radical230.3

OpenStudy (jdoe0001):

\(\bf \sqrt{7.6}+\sqrt{230.3}\quad ?\)

OpenStudy (anonymous):

Yes but a 2 in front of both radicals

OpenStudy (jdoe0001):

\(\bf 2\sqrt{7.6}+2\sqrt{230.3}\) hmm I'd think you could just use your calculator

OpenStudy (anonymous):

Ok thanks